Mark, Andrew and Dick shared some money. Mark had 480 dollars more than Andrew. Dick had 3 times as much as Mark. When Dick used 600 dollars, the total amount of money that Dick and Andrew had was 1500 dollars.
How much did each one have.
Thanks

Answer by richwmiller(17219)

m=a+480
d=3*m
d-600+a=1500
substitute
3*m-600+a=1500
substitute again
3*(a+480)-600+a=1500
3*a+3*480-600+a=1500
3*a+3*480-600+a=1500
4*a=1500+600-1440
4*a=660
a=660/4
a=165 for Andrew
m=a+480
m=165+480=645 for Mark
d=3*m
d=3*645=1935 for Dick
check
1935-600=1335
165+1335=1500
ok